OpenStudy (anonymous):

What is the integral of ln(x+1) on (1,2)?

sam (.sam.):

Use int by parts

OpenStudy (anonymous):

is that the correct integral?\[\int\limits_{1}^{2} \ln(x+1) dx\]

OpenStudy (anonymous):

I tried, but it still didn't come out correctly... I also tried making u=x+1 and then using int by parts, but the answer the lesson gave me was also different (xln(x+1) - x + ln(x+1) on (1,2)). lol so i have no idea how they got that

OpenStudy (anonymous):

yes, that's the right integral :)

sam (.sam.):

\[\int\limits \ln (x+1) \, dx\] u=x+1 du=dx ------------------------------------------------------------- \[\int\limits \ln u \, du\] Int by parts \[\int\limits f \, dg=f g-\int\limits g \, df\] \[f=\ln(u)~~~~~~~dg=du \\ \\ df=\frac{1}{u} du~~~~~~~g=u\] ------------------------------------------------------------- \[u (\ln(u) )-\int\limits 1 \, du\] \[uln(u)-u+c\]

OpenStudy (anonymous):

ok, i got the same thing you did, which actually turns out to be (x+1)ln(x+1) - x - 1 instead of xln(x+1) -x+ln(x+1), so I guess they just got it wrong or something. Thanks!

sam (.sam.):

Are (1,2) coordinates?

OpenStudy (anonymous):

no, they're the limits of integration.

sam (.sam.):

Okay \[\Large (x+1)(\ln(x+1))-(x+1)|_1^2\]

OpenStudy (anonymous):

ok, so I plugged that into my calculator, but the answers were still different: -.1925 as opposed to 0.9095

sam (.sam.):

You should get 0.9095 did you typed it correctly?

OpenStudy (anonymous):

I thought so, but maybe not. I used fnInt((x+1)(ln(x+1)) - (x+1), x, 1, 2)?

sam (.sam.):

(2+1)(ln(2+1))-(2+1)-[(1+1)(ln(1+1)-(1+1)]

OpenStudy (anonymous):

oh never mind! haha i didn't use the integral

OpenStudy (anonymous):

ok, I got it. Thanks!

sam (.sam.):

welcome :)